OpenStudy (anonymous):

The amount of radioactive uranium changes with time. The table below shows the amount of radioactive uranium f(t) left after time t: t(hours) 0 0.5 1 f(t) 100 50 25 Which exponential function best represents the relationship between f(t) and t?

OpenStudy (anonymous):

@mathstudent55 @hba

OpenStudy (haseeb96):

at every half of increase of time the amount of radioactive material is also half

OpenStudy (anonymous):

f(t) = 100(0.25)t f(t) = 0.25(100)t f(t) = 100 (0.5)t f(t) = 0.25(50)t

OpenStudy (anonymous):

those are the chocies ^^^

OpenStudy (mathstudent55):

Look at you choices. Insert each given point in all your choices. The choice that works for all points is the answer.

OpenStudy (anonymous):

I dont understand how to do it

OpenStudy (anonymous):

oh wait is it d

OpenStudy (mathstudent55):

Point 1 is (0, 100) Let's try point 1 in choice 1: f(t) = 100(0.25)t f(0) = 100(0.25)^0 = 100 * 1 = 100 Choice 1 works for point 1

OpenStudy (mathstudent55):

Now we need to test point 2 in choice 1.

OpenStudy (mathstudent55):

Point 2 is (0.5, 50) Choice 1 is f(t) = 100(0.25)t so we get f(0.5) = 100(0.25)^0.5 = 100 * 0.5 = 50 It works.

OpenStudy (mathstudent55):

Let's try point 3 with choice 1. Point 3 is (1, 25) f(t) = 100(0.25)t f(1) = 100(0.25)^1 = 100 * 0.25 = 25 It also works. This is the answer.

OpenStudy (anonymous):

thanks!! so much!!!

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

what was the anwser?